# Pipeline Parallelism with Controllable Memory

### v3

@@ 3.4 Other Building Blocks
Besides V-Shape building blocks, we also propose some other interesting building blocks in Appendix H, to show the generalization ability of our framework. Some useful examples include a) 1F1B-V achieving 2/3 of 1F1B’s activation memory without doing B-W split; b) a schedule with the same bubble rate as Interleaved 1F1B(Narayanan et al., 2021) but lower memory consumption. Additionally, we design an adaptive scheduler to control the memory at a finer granularity in Appendix A.
@@ 4 Experiments
We construct our experiments to show three conclusions: a) The throughput and memory of V-Min, V-Half and V-Min aligns with the theoretical analysis in Section 3.2; b) Memory-saving methods including V-Min and V-ZB can bring accelerations; c) Our methods still perform best when combining with other state-of-the-art techniques.
@@ 4.1 Setup
We evaluate our methods using a serious of models detailed in Table 2 analogous to GPT-3 (Brown et al., 2020). Our implementation is based on open-source Megatron-LM project (Narayanan et al., 2021) and is experimented on up to 40 NVIDIA A100 SXM 80G GPUs distributed across 4 nodes interconnected by a RoCE RDMA network. The running time of each iteration is recorded after several warm-up iterations. Similar to the settings in (Qi et al., 2023), we deduct one transformer layer from both the initial and final pipeline stage to compensate for the embedding and output layer in LM, which can otherwise become the bottleneck of the pipeline and interfere the efficiency.
@@
Figure 6: Throughput and activation memory using same micro batch size.
@@ 4.2 Comparing Pipeline Schedules
In Figure 6, we present comparisons of the throughput measured in FLOPS utilization (MFU) and activation memory consumption across different pipeline schedules under various settings. From the results, V-ZB outperforms all other methods in terms of throughput, which aligns with Figure 4. When comparing the activation memory consumption, V-Min and V-Half stand out by significantly reducing activation memory to approximately 1/3 and 1/2, while other methods’ memory is similar. More details of our experiments and definition of metrics can be found in Appendix D.1.
@@
Figure 7: Pareto Frontier of MFU and memory for various setup.
@@ 4.4 Combining with Existing Techniques
We present the best result for each pipeline parallel schedule in Table 3 and the corresponding parameters. We find that when sequence length is smaller and hence the memory budget is more abundant, V-ZB performs the best due to the elimination of bubbles. When we increase the sequence length and the number of microbatches, V-Half performs the best because of its memory efficiency. The detailed data and analysis of grid search can be found in the Appendix D.3.
Table 3: V Schedules Combined with Other Memory Saving Methods
@@
1F1B 49.81 2 4 5 2
98.5B Model 1F1B-R 37.15 2 2 10 1
Seq Length 1024 ZB-1P 54.52 1 1 40 1
Batch Size 160 V-Half 52.60 2 1 20 1
V-Min 48.02 2 2 10 1
V-ZB 57.58 1 1 40 1
1F1B 57.25 2 4 5 1
98.5B Model 1F1B-R 45.81 2 1 20 1
Seq Length 3072 ZB-1P 56.55 1 4 10 1
Batch Size 640 V-Half 60.34 1 2 20 1
V-Min 55.51 1 1 40 1
V-ZB 56.89 1 4 10 1
@@ 5 Conclusion And Future Work
In this work, we present a framework that constructs pipeline schedules by focusing on their repeating building blocks. This framework enables direct computation of peak memory from the lifespan of the building block. Based on this capability, we design a family of memory-efficient building blocks. We discuss three representative methods from this family, namely V-Min, V-Half and V-ZB, and demonstrate with experiments that our methods advance the pareto frontier of throughput and memory in large model training. Furthermore, our methodology of designing pipeline schedules through building blocks may inspire the research community to explore more novel pipeline schedules.
@@ Appendix B Evaluation of Adaptive Scheduler
In Figure 9 we plot the bubble rate of adaptive V-Shape schedulers introduced in A under various settings and different memory limit. The run times of F, B and W are from profiled real cases, as in Table 5. We observe that the bubble rate drops as memory limit increases. Notably, there’s a sudden drop of bubble rate when the memory limit just goes above approximately 1/2 of 1F1B, at which point the per-microbatch bubble mentioned in Figure 5(a) disappears.
@@ D.1 Comparing Pipeline Schedules
For Section 4.2, we present our detailed experiment data in Table 4. Specifically the metrics are defined as:
@@ D.3 More Details on Grid Search
Parallelisation MicroBS 1F1B 1F1B-R ZB-1P V-Half V-Min V-ZB
DP=1 1 46.99 34.72 47.83 47.59 43.07 48.58
TP=4 2 49.11 36.67 51.27 50.86 45.69 52.98
PP=10 4 47.69 35.3 - 50.46 45.29 -
DP=2 1 49.26 37.15 52.08 51.7 48.02 53.69
TP=2 2 - 36.7 - 51.87 47.56 -
PP=10 4 - 31.96 - - - -
DP=1 1 49.14 36.88 52.43 51.73 46.02 54.19
TP=2 2 48.47 36.22 52.34 51.75 45.84 54.78
PP=20 4 - 31.48 - - 41.4 -
DP=2 1 - 36.69 - 52.6 48.0 -
TP=1 2 - 32.64 - - 44.41 -
PP=20 4 - 25.27 - - - -
DP=1 1 48.63 36.44 54.52 52.46 47.78 57.58
TP=1 2 - 32.46 - 49.19 43.91 -
PP=40 4 - 25.04 - - - -
DP=1 1 40.66 30.2 40.56 40.07 33.54 41.22
TP=8 2 47.07 35.17 47.9 47.24 41.26 48.45
PP=5 4 47.3 35.01 48.96 48.19 42.45 49.09
8 45.66 33.5 - 47.17 42.59 -
DP=2 1 46.89 35.22 48.55 47.79 44.3 48.75
TP=4 2 49.81 37.01 52.27 51.22 47.51 52.79
PP=5 4 - 36.03 - 51.41 47.43 -
We show the MFU of every setup of our grid search in Table 6 and 7 for two groups of experiments: one with $SequenceLength=1024$ and $BatchSize=160$, and the other with $SequenceLength=3072$ and $BatchSize=640$.
For the first experiment group, the best setup is V-ZB under pure PP because of its bubble elimination. For the second setup, the best setup is V-Half because its memory efficiency enables a lower TP degree which is otherwise impossible for V-ZB/ZB-1P/1F1B. A comparison of TP and PP can be found at D.4.
@@
Parallelisation MicroBS 1F1B 1F1B-R ZB-1P V-Half V-Min V-ZB
d=1 1 56.45 41.7 56.55 56.54 51.19 56.89
t=4 2 - 41.4 - 41.71 50.52 -
p=10 4 - 41.24 - - - -
d=2 1 - 44.59 - - 55.11 -
t=2 2 - 43.1 - - - -
p=10 4 - 41.25 - - - -
d=1 1 - 44.23 - 60.34 52.68 -
t=2 2 - 42.91 - - - -
p=20 4 - 41.2 - - - -
d=2 1 - 45.81 - - - -
t=1 2 - 41.85 - - - -
p=20 4 - - - - - -
d=1 1 - 45.2 - - 55.51 -
t=1 2 - 41.16 - - - -
p=40 4 - 38.86 - - - -
d=1 1 50.87 37.57 49.23 49.31 45.25 49.46
t=8 2 52.42 38.54 50.25 50.6 47.18 50.71
p=5 4 - 39.38 - - 13.27 -
d=2 1 57.25 42.28 - 56.35 52.51 -
t=4 2 - 41.76 - - - -
p=5 4 - 40.42 - - - -
@@ D.4 Model Parallelism: More PP or More TP?
The throughput of PP also decreases as PP degree increases for two major reasons: a) for PP schedules with a fixed global batch size, a higher PP degree usually results in higher pipeline bubbles. For example, 1F1B has a bubble rate of $\frac{p-1}{p+n-1}$, which would increase if $p$ grows and $n$ keeps unchanged; b) the first and last pipeline stage for a language model usually have an embedding or output layer which has innegligible additional computations, making the pipeline unbalanced and hurting the efficiency. With higher PP degree, the unbalance would be aggravated. However theses two issues can be mitigated. For the first issue, a larger number of microbatches or using V-ZB can significantly reduce the pipeline bubble. For the second issue, we can use another standalone stage to handle the embedding or output layer instead to avoid bottleneck. In our experiments, we simply deduct 1 layer from both the first and last pipeline stage to mitigate this problem. As a result, our methods essentially push the preference from TP to PP.
@@ E.1 Existence of Repeating Bubbles
If there is no repeating bubbles in a schedule for any values of $T_{F},T_{B},T_{W}$, the minimal memory is $2dm$.
@@
Then we consider a similar dependency path $\tau_{3}$ containing only backward passes and we can get $\delta(B_{1}^{1},B_{1}^{0})>6d-6$. According to Equation 1, we can get the peak memory in the first device is at least $(\lceil\frac{\delta(F_{1}^{0},F_{1}^{1})}{6}\rceil+\lceil\frac{\delta(B_{1}^{1},B_{1}^{0})}{6}\rceil)m\geq 2dm.$
For most real cases, V-Half is enough to guarantee there is no repeating bubbles.
@@
Because both warm-up phase and cool-down phase have $O(d)$ passes, we only need to consider the stable phase to identify whether a schedule is with $O(d)$ bubble or with $O(n)$ bubble. Obviously, streaming dependency won’t block the device from executing next pass. So bubble is caused only by the computational dependency. Formally, a schedule is with $O(n)$ bubble if and only if there exist two passes $u$ and $v$ (within the same device) and there are two dependency paths $\tau$ and $\tau^{\prime}$ between them, where $\tau$ only contains streaming dependencies, $\tau^{\prime}$ contains at least two computational dependencies, and $T_{\tau}<T_{\tau^{\prime}}$. Based on our V-shape building blocks, we only need to check $u$ and $v$ with a small distance ($<6$) and $\tau^{\prime}$ within two adjacent devices. In this way, we can conclude that the schedule in Figure 4(c) is with $O(d)$ bubble when $T_{W}+2T_{B}\geq 2T_{F}\And T_{W}+2T_{F}\geq 2T_{B}$, which is satisfied in most real cases.
@@ Appendix F Building blocks of V-Min and V-Half for all values of $d$
To avoid collisions when repeating building blocks, we need a slightly different offsets for different number of devices $d$. We list the details in Table 8 and 9 while continue using the notation defined in A. We do not list the offsets related to W passes because they always fill in the empty grids left by F and B. Some samples can also be found at Figure 14.
@@ Appendix G Non-uniform Repeating Interval of Interleaved 1F1B
Appendix H Other Memory Efficient Building Blocks
Under the guidance of lifespan defined in Section 2.2, we can also find some other building blocks besides the V-Shape building block family. We show the building blocks in Figure 16 and their full schedules in Figure 17. All these schedules have lower bubble rate than 1F1B.
Specifically, 1F1B-V applies V-Shape to the building block of 1F1B but without B-W split, which can reduce the peak memory to asymptotically 2/3 of 1F1B.
We also find that utilizing B-W split, the zero bubble pipeline schedules proposed in [Qi et al., 2023] with configurable memory limit can support a minimum of 2/3 activation memory of 1F1B, using the building block shown in Figure 16(b). Note that two microbatches are included in a single building block to avoid collision.
Using the building block defined in Figure 16(c), we can make a schedule with the same bubble rate as interleaved 1F1B but lower memory, shown in Figure 17(c).
@@ Appendix I A Gallery of Pipeline Parallel Schedules and Their Building Blocks
(i) Interleaved 1F1B with uniform interval
Figure 18: A gallery of pipeline schedules and their building blocks. The upper row of each schedule shows the building block and how it repeats. The lower row shows the final schedule after squeezing and reordering.

### v4

@@ 3.4 Other Building Blocks
Besides V-Shape building blocks, we also propose some other interesting building blocks in Appendix H, to show the generalization ability of our framework. Some useful examples include a) 1F1B-V achieving 2/3 of 1F1B’s activation memory without doing B-W split; b) a schedule consumes less memory than interleaved 1F1B but with the same bubble rate (Figure 17(c)). Additionally, we design an adaptive scheduler to control the memory at a finer granularity in Appendix A.
@@ 4 Experiments
We construct our experiments to show three conclusions: a) The throughput and memory of V-Min, V-Half and V-ZB aligns with the theoretical analysis in Section 3.2; b) Memory-saving methods including V-Min and V-Half can bring accelerations; c) Our methods still perform best when combining with other state-of-the-art techniques.
@@ 4.1 Setup
We evaluate our methods using a series of models detailed in Table 2 analogous to GPT-3 (Brown et al., 2020). Our implementation is based on the open-source Megatron-LM project (Narayanan et al., 2021) and is experimented on up to 40 NVIDIA A100 SXM 80G GPUs distributed across 5 nodes interconnected by a RoCE RDMA network. The running time of each iteration is recorded after several warm-up iterations. Similar to the settings in (Qi et al., 2023), we deduct one transformer layer from both the initial and final pipeline stage to compensate for the embedding and output layer in LM, which can otherwise become the bottleneck of the pipeline and interfere to the efficiency.
@@
Figure 6: Throughput and activation memory using the same microbatch size.
@@ 4.2 Comparing Pipeline Schedules
In Figure 6, we present comparisons of the throughput measured in FLOPS utilization (MFU) and activation memory consumption across different pipeline schedules under various settings. From the results, V-ZB outperforms all other methods in terms of throughput, which aligns with Figure 4. When comparing the activation memory consumption, V-Min and V-Half stand out by significantly reducing activation memory to approximately 1/3 and 1/2, while other methods’ memory is similar except for 1F1B-R. More details of our experiments and definition of metrics can be found in Appendix D.1.
@@
Figure 7: Pareto frontier of MFU and memory for various setups.
@@ 4.4 Combining with Existing Techniques
We present the best result for each pipeline parallel schedule in Table 3 and the corresponding parameters. We find that when sequence length is smaller and hence the memory budget is more abundant, V-ZB performs the best due to the elimination of bubbles. When we increase the memory pressure by increasing the sequence length, V-Half performs the best because of its memory efficiency. The detailed data and analysis of grid search can be found in the Appendix D.3 D.4.
Table 3: V-Shape schedules combined with other memory saving methods.
@@
1F1B 54.77 2 4 5 2
98.5B Model 1F1B-R 40.84 2 2 10 1
Seq Length 1024 ZB-1P 59.95 1 1 40 1
Batch Size 160 V-Half 57.83 2 1 20 1
V-Min 52.8 2 2 10 1
V-ZB 63.31 1 1 40 1
1F1B 62.95 2 4 5 1
98.5B Model 1F1B-R 50.37 2 1 20 1
Seq Length 3072 ZB-1P 62.18 1 4 10 1
Batch Size 640 V-Half 66.34 1 2 20 1
V-Min 61.04 1 1 40 1
V-ZB 62.56 1 4 10 1
1F1B OOM -
98.5B Model 1F1B-R 42.05 1 4 10 1
Seq Length 16384 ZB-1P OOM -
Batch Size 160 V-Half 57.85 1 8 5 1
V-Min 48.58 1 8 5 1
V-ZB OOM -
@@ 5 Conclusion And Future Work
In this work, we present a framework that constructs pipeline schedules by focusing on their repeating building blocks. This framework enables direct computation of peak memory from the lifespan of the building block. Based on this capability, we design a family of memory-efficient building blocks. We discuss three representative methods from this family, namely V-Min, V-Half and V-ZB, and demonstrate with experiments that our methods advance the Pareto frontier of throughput and memory in large model training. Furthermore, our methodology of designing pipeline schedules through building blocks may inspire the research community to explore more novel pipeline schedules. Notice that repeating a building block is not the only way of building a pipeline, other methods like greedy search could generate a pipeline that has no repeating patterns.
@@ Appendix B Evaluation of Adaptive Scheduler
In Figure 9 we plot the bubble rate of adaptive V-Shape schedulers introduced in A under various settings and different memory limit. The run times of F, B and W are from profiled real cases, as in Table 5. We observe that the bubble rate drops as memory limit increases. Notably, there’s a sudden drop in bubble rate when the memory limit just goes above approximately 1/2 of 1F1B, at which point the repeating bubble mentioned in Figure 5(a) disappears.
@@ D.1 Comparing Pipeline Schedules
For Section 4.2, we present our detailed experiment data in Table 4. Specifically, the metrics are defined as:
@@ D.3 More Details on Grid Search
Parallelization MicroBS 1F1B 1F1B-R ZB-1P V-Half V-Min V-ZB
DP=1 1 51.66 38.18 52.59 52.32 47.36 53.42
TP=4 2 54.0 40.32 56.37 55.92 50.24 58.25
PP=10 4 52.44 38.81 - 55.49 49.79 -
DP=2 1 54.17 40.84 57.26 56.85 52.8 59.03
TP=2 2 - 40.35 - 57.03 52.3 -
PP=10 4 - 35.14 - - - -
DP=1 1 54.03 40.55 57.65 56.88 50.61 59.59
TP=2 2 53.3 39.83 57.55 56.9 50.4 60.23
PP=20 4 - 34.62 - - 45.52 -
DP=2 1 - 40.34 - 57.83 52.78 -
TP=1 2 - 35.89 - - 48.83 -
PP=20 4 - 27.78 - - - -
DP=1 1 53.47 40.07 59.95 57.68 52.54 63.31
TP=1 2 - 35.69 - 54.08 48.27 -
PP=40 4 - 27.53 - - - -
DP=1 1 44.71 33.21 44.6 44.05 36.88 45.32
TP=8 2 51.75 38.67 52.67 51.94 45.37 53.27
PP=5 4 52.0 38.49 53.84 52.98 46.67 53.98
8 50.2 36.84 - 51.87 46.83 -
DP=2 1 51.56 38.73 53.38 52.54 48.7 53.6
TP=4 2 54.77 40.69 57.48 56.32 52.24 58.05
PP=5 4 - 39.61 - 56.53 52.16 -
We show the MFU of every setup of our grid search in Table 6, 7 and 8 for three groups of experiments: one with $SequenceLength=1024$ and $BatchSize=160$, one with $SequenceLength=3072$ and $BatchSize=640$ and the other with $SequenceLength=16384$ and $BatchSize=160$.
For the first experiment group, the best setup is V-ZB under pure PP because of its bubble elimination. For the second setup, the best setup is V-Half because its memory efficiency enables a lower TP degree, which is otherwise impossible for V-ZB/ZB-1P/1F1B. For the last setup, due to high memory pressure only V-Min and V-Half can run without checkpointing. A comparison of TP and PP can be found at D.4.
@@
Parallelization MicroBS 1F1B 1F1B-R ZB-1P V-Half V-Min V-ZB
DP=1 1 62.06 45.85 62.18 62.17 56.29 62.56
TP=4 2 - 45.52 - 45.86 55.54 -
PP=10 4 - 45.34 - - - -
DP=2 1 - 49.03 - - 60.59 -
TP=2 2 - 47.39 - - - -
PP=10 4 - 45.35 - - - -
DP=1 1 - 48.63 - 66.34 57.92 -
TP=2 2 - 47.18 - - - -
PP=20 4 - 45.3 - - - -
DP=2 1 - 50.37 - - - -
TP=1 2 - 46.01 - - - -
PP=20 4 - - - - - -
DP=1 1 - 49.7 - - 61.04 -
TP=1 2 - 45.26 - - - -
PP=40 4 - 42.73 - - - -
DP=1 1 55.93 41.31 54.13 54.22 49.75 54.38
TP=8 2 57.64 42.37 55.25 55.64 51.88 55.76
PP=5 4 - 43.3 - - 14.59 -
DP=2 1 62.95 46.49 - 61.95 57.74 -
TP=4 2 - 45.92 - - - -
PP=5 4 - $4$
Table 8: MFU of grid search, with $SequenceLength=16384$ and $BatchSize=160$
Parallelization MicroBS 1F1B 1F1B-R ZB-1P V-Half V-Min V-ZB
DP=1;TP=4;PP=10 1 - 42.05 - - - -
DP=2;TP=2;PP=10 1 - - - - - -
DP=1;TP=2;PP=20 1 - 41.52 - - - -
DP=2;TP=1;PP=20 1 - - - - - -
DP=1;TP=1;PP=40 1 - - - - - -
DP=1;TP=8;PP=5 1 - 39.48 - 57.85 48.58 -
DP=1;TP=8;PP=5 2 - - - - - -
DP=2;TP=4;PP=5 1 - 35.13 - - - -
@@ D.4 Model Parallelism: More PP or More TP?
The throughput of PP also decreases as PP degree increases for two major reasons: a) for PP schedules with a fixed global batch size, a higher PP degree usually results in higher pipeline bubbles. For example, 1F1B has a bubble rate of $\frac{p-1}{p+n-1}$, which would increase if $p$ grows and $n$ keeps unchanged; b) the first and last pipeline stage for a language model usually have an embedding or output layer which has innegligible additional computations, making the pipeline unbalanced and hurting the efficiency. With higher PP degree, the unbalance would be aggravated. However, theses two issues can be mitigated. For the first issue, a larger number of microbatches or using V-ZB can significantly reduce the pipeline bubble. For the second issue, we can use another standalone stage to handle the embedding or output layer instead to avoid bottleneck. In our experiments, we simply deduct 1 layer from both the first and last pipeline stage to mitigate this problem. As a result, our methods essentially push the preference from TP to PP.
@@ E.1 Existence of Repeating Bubbles
If there are no repeating bubbles in a schedule for any values of $T_{F},T_{B},T_{W}$, the minimal memory is $2dm$.
@@
Then we consider a similar dependency path $\tau_{3}$ containing only backward passes, and we can get $\delta(B_{1}^{1},B_{1}^{0})>6d-6$. According to Equation 1, we can get the peak memory in the first device is at least $(\lceil\frac{\delta(F_{1}^{0},F_{1}^{1})}{6}\rceil+\lceil\frac{\delta(B_{1}^{1},B_{1}^{0})}{6}\rceil)m\geq 2dm.$
For most real cases, V-Half is enough to guarantee there are no repeating bubbles.
@@
Because both warm-up phase and cool-down phase have $O(d)$ passes, we only need to consider the stable phase to identify whether a schedule is with $O(d)$ bubble or with $O(n)$ bubble. Obviously, streaming dependency won’t block the device from executing the next pass. So bubble is caused only by the computational dependency. Formally, a schedule is with $O(n)$ bubble if and only if there exist two passes $u$ and $v$ (within the same device) and there are two dependency paths $\tau$ and $\tau^{\prime}$ between them, where $\tau$ only contains streaming dependencies, $\tau^{\prime}$ contains at least two computational dependencies, and $T_{\tau}<T_{\tau^{\prime}}$. Based on our V-shape building blocks, we only need to check $u$ and $v$ with a small distance ($<6$) and $\tau^{\prime}$ within two adjacent devices. In this way, we can conclude that the schedule in Figure 4(c) is with $O(d)$ bubble when $T_{W}+2T_{B}\geq 2T_{F}\And T_{W}+2T_{F}\geq 2T_{B}$, which is satisfied in most real cases.
@@ Appendix F Building blocks of V-Min and V-Half for all values of $d$
To avoid collisions when repeating building blocks, we need a slightly different offsets for different number of devices $d$. We list the details in Table 9 and 10 while continue using the notation defined in A. We do not list the offsets related to W passes because they always fill in the empty grids left by F and B. Some samples can also be found at Figure 14.
@@ Appendix G Non-uniform Repeating Interval of Interleaved 1F1B
Appendix H Other Memory Efficient Building Blocks
Under the guidance of lifespan defined in Section 2.2, we also find some other building blocks besides the V-Shape building block family. We show the building blocks in Figure 16 and their full schedules in Figure 17. All these schedules have lower bubble rate than 1F1B. Specifically, 1F1B-V applies V-Shape to the building block of 1F1B but without B-W split, which can reduce the peak memory to asymptotically 2/3 of 1F1B. We also find that utilizing B-W split, the zero bubble pipeline schedules proposed in [Qi et al., 2023] with configurable memory limit can support a minimum of 2/3 activation memory of 1F1B, using the building block shown in Figure 16(b). Note that two microbatches are included in a single building block to avoid collision. Using the building block defined in Figure 16(c), we can make a schedule with the same bubble rate as interleaved 1F1B but lower memory, shown in Figure 17(c).
@@ Appendix I A Gallery of Pipeline Parallel Schedules and Their Building Blocks
(i) Interleaved 1F1B with uniform interval (with reordering)
Figure 18: A gallery of pipeline schedules and their building blocks. The upper row of each schedule shows the building block and how it repeats. The lower row shows the final schedule after squeezing.